# Continuous-Aperture Array (CAPA)-Based Wireless Communications: Capacity Characterization

### v1

@@ III Single-User Case
III-A Uplink Capacity
We commence with the uplink scenario. In the single-user case, (3b) simplifies to
${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})=\mathsf{H}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}}}A_{\mathsf{u}}s_{{\mathsf{ul}}}+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}).$ (11)
We then design a detector ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$ to recover $s_{{\mathsf{ul}}}$ from ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$. Specifically, applying this detector to ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$ yields
@@ III-A Uplink Capacity
The signal-to-noise ratio (SNR) for decoding $s_{{\mathsf{ul}}}$ is thus given by
@@
The subsequent task is to design ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$ to maximize $\gamma_{\mathsf{ul}}$. As observed from (13), $\gamma_{\mathsf{ul}}$ is independent of the norm of ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$, i.e., $\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}$. Thus, the optimal detector satisfies
@@
The results in (14) suggest that the capacity-achieving detector for uplink single-user CAPA communications aligns with the spatial channel response ${\mathsf{H}}(\mathbf{r})$, which can be regarded as the continuous version of maximal-ratio combining (MRC).
Substituting (14) back into (13) gives
@@
III-B Downlink Capacity
Turning to the downlink case, for $K=1$, (7b) simplifies to
where ${\mathbbmss{E}}\{{\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}^{*}({\mathbf{s}}^{\prime})\}={\sigma}_{\mathsf{u}}^{2}\delta(\mathbf{s}-{\mathbf{s}}^{\prime})$. Applying the detector ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$ to ${\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s})$ yields
By noting that ${A_{\mathsf{u}}}$ is negligible with respect to the propagation distance, we have
$\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\approx{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}_{\mathsf{u}})A_{\mathsf{u}},\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}}\approx\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}},$
which simplifies (19) as follows:
$\gamma_{\mathsf{dl}}$ $\approx\frac{\lvert{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}}^{2}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}}}$ (20a)
$=\frac{A_{\mathsf{u}}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}}.$ (20b)
The results in (20b) suggest that the downlink SNR is independent of the detector ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$ under our settings. Thus, any detector can be used to recover $s_{{\mathsf{dl}}}$. Additionally, from (20b), it is observed that the SNR is maximized when the current distribution satisfies ${\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\propto{\mathsf{H}}^{*}({\mathbf{r}})$, i.e.,
Remark 2.
Inserting (21) into (20b) gives
IV Multiuser Case
This section extends the single-user scenario to the multiuser one. We characterize the channel capacity for both uplink and downlink communications. For brevity, our analysis focuses on the two-user case, while extensions to scenarios with an arbitrary number of users will be addressed in future work.
IV-A Uplink Capacity
We first analyze the uplink scenario.
IV-A1 Optimal Uplink Detection
@@
which is the same as that given in (11). Following the derivation steps to obtain (14), the BS can use the detector ${\mathsf{V}}_{{\mathsf{ul}},1}(\mathbf{r})=\frac{{\mathsf{H}_{1}}({\mathbf{r}})}{\sqrt{\int_{{\mathcal{A}}}\lvert{\mathsf{H}}_{1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}}}$ to recover $s_{{\mathsf{ul}},1}$ from (25). Therefore, the SNR for decoding the message sent by user $1$ is given by
@@
    Whitening: The autocorrelation function of ${\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Z}}(\mathbf{r}){\rm{d}}\mathbf{r}$ is proportional to the Dirac delta function, i.e.,
@@
Next, we exploit ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ to transform ${\mathsf{Y}}_{\mathsf{B}}(\mathbf{r})$ as follows:
@@
where $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}\mathbf{r}$. By noting that ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})$, the model presented in (31) is similar to that in (11). Motivated by this, we design an MRC-based detector $\mathsf{V}_{\mathsf{ul},2}(\mathbf{r}^{\prime})$ by treating $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})$ as the equivalent channel response, which yields
@@ IV-A2 Capacity Characterization
IV-B Downlink Capacity
In this part, we focus on the downlink scenario.
@@ IV-B1 Optimal Multiuser Precoding
where ${\mathsf{v}}_{\varepsilon_{k}}=\int_{\mathcal{A}_{\varepsilon_{k}}}{\lvert\mathsf{V}_{{\mathsf{dl}},\varepsilon_{k}}(\mathbf{s})\rvert^{2}}\mathrm{d}\mathbf{s}$ and $\overline{\mathsf{v}}_{\varepsilon_{k}}=\lvert\int_{\mathcal{A}_{\varepsilon_{k}}}{\mathsf{V}}_{{\mathsf{dl}},\varepsilon_{k}}^{*}(\mathbf{s})\mathrm{d}\mathbf{s}\rvert^{2}$. Since each user’s array aperture is negligible with respect to the propagation distance, we have
@@
where $\mathsf{P}$ reflects the transmit power. Following the same derivation approach, the sum-rate achieved by the DPC order $1\rightarrow 2$ can also be derived. For Gaussian channels, the maximal DPC sum-rate is independent of the DPC decoding order, and thus $\mathsf{C}_{\mathsf{dl}}^{2\rightarrow 1}$ is essentially the downlink sum-rate capacity [21].
@@
$\begin{split}{\mathsf{C}}_{\mathsf{du}}=\max_{{\mathsf{P}}_{1},{\mathsf{P}}_{2}\geq 0,{\mathsf{P}}_{1}+{\mathsf{P}}_{2}\leq{\mathsf{P}}}&\log_{2}(1+\overline{\gamma}_{\mathsf{du},1}\mathsf{g}_{1}+\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{2}\\ &+\overline{\gamma}_{\mathsf{du},1}\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{1}\mathsf{g}_{2}\overline{\rho}),\end{split}$ (50)
@@
Having obtained the capacity-achieving source currents under the DPC order $2\rightarrow 1$, we summarize the entire precoding procedure in Table II for ease of reference. The presented DPC procedure can be directly extended to the other order $1\rightarrow 2$ by exchanging the user indices.
@@
TABLE II: DPC-based precoding for downlink CAPA communications.
@@ IV-B2 Capacity Characterization
As proved in Appendix -D, given the power allocation policy $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$, the achievable rates of user $1$ and user $2$ under the DPC order $2\rightarrow 1$ are given as follows:
@@
$\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1}$ $=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}\right).$ (55)
By comparing the above two equations with the results in Corollary 1, the following observations are made.
@@
The achievable rates under the DPC order $1\rightarrow 2$ can be derived with similar steps, which are omitted for brevity. For a thorough study, we derive the downlink-to-uplink transformation as follows.
@@
=$=\begin{cases}\log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},1}(\mathsf{P})\mathsf{g}_{1}\right)&\xi\geq{\mathsf{P}}\\ \log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},2}(\mathsf{P})\mathsf{g}_{2}\right)&\xi\leq-{\mathsf{P}}\\ \log_{2}\left(1+\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{1}\varepsilon_{2}\overline{\rho}\right)&\mathrm{else}\\ \end{cases},$ (58b)
where $\widetilde{\gamma}_{\mathsf{dl},k}(x)\triangleq\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}x$ for $k=1,2$, $\varepsilon_{1}=\widetilde{\gamma}_{\mathsf{dl},1}(\frac{{\mathsf{P}}-\xi}{2})\mathsf{g}_{1}$, and $\varepsilon_{2}=\widetilde{\gamma}_{\mathsf{dl},2}(\frac{{\mathsf{P}}+\xi}{2})\mathsf{g}_{2}$.
@@
Furthermore, according to the uplink-downlink duality, given a power allocation scheme $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$ with ${\mathsf{P}}_{1}+{\mathsf{P}}_{2}={\mathsf{P}}$, we can obtain the capacity region of a dual uplink channel. The downlink capacity region is the convex hull of the union of all these dual uplink capacity regions, which is illustrated in Section VI.

### v2

@@ III-A Single-User Case
The signal-to-noise ratio (SNR) for decoding $s_{{\mathsf{ul}}}$ reads
@@
It can be observed from (12) that the optimal detector satisfies
@@
The results in (13) suggest that the capacity-achieving detector for uplink single-user CAPA communications aligns with the spatial channel response ${\mathsf{H}}(\mathbf{r})$, which is the continuous version of maximal-ratio combining (MRC).
Substituting (13) into (12) gives
@@
III-B Two-User Case
We now extend the single-user scenario to the two-user one.
III-B1 Optimal Uplink Detection
@@
Following the derivation steps to obtain (13), the BS can use the detector ${\mathsf{V}}_{{\mathsf{ul}},1}(\mathbf{r})=\frac{{\mathsf{H}_{1}}({\mathbf{r}})}{\sqrt{\int_{{\mathcal{A}}}\lvert{\mathsf{H}}_{1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}}}$ to recover $s_{{\mathsf{ul}},1}$ from (17). Therefore, the SNR for decoding the message sent by user $1$ is given by
@@
    Whitening: The autocorrelation function of ${\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Z}}(\mathbf{r}){\rm{d}}\mathbf{r}$ aligns with the delta function, i.e.,
@@
Next, we exploit ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ to transform ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$ as follows:
@@
where $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}\mathbf{r}$. By noting that ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})$, the model presented in (23) is similar to that in the single-user case. Motivated by this, we design an MRC-based detector $\mathsf{V}_{\mathsf{ul},2}(\mathbf{r}^{\prime})$ by treating $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})$ as the equivalent channel response, which yields
@@ III-B2 Capacity Characterization
III-B3 Sum-Rate Achieved by Zero-Forcing Detector
To attain the uplink sum-rate capacity in (33), a non-linear detector based on SIC is necessary. However, in practical systems, the linear ZF detector is frequently favored to lower decoding complexity. In the subsequent analysis, we aim to evaluate the sum-rate performance of the ZF detector and contrast it with the sum-rate capacity achieved by the non-linear detector.
Let $\mathsf{V}_{\mathsf{zf},k}(\mathbf{r})$ denote the ZF detector for user $k$ with $k\in\{1,2\}$. In order to nullify the IUI, i.e., $\int_{{\mathcal{A}}}\mathsf{V}_{\mathsf{zf},k}^{*}(\mathbf{r})\mathsf{H}_{k^{\prime}}(\mathbf{r}){\rm{d}}{\mathbf{r}}=0$ for $k^{\prime}\neq k$, and maximize the channel gain of user $k$, $\mathsf{V}_{\mathsf{zf},k}(\mathbf{r})$ is designed as the projection of $\mathsf{H}_{k}(\mathbf{r})$ onto the space orthogonal to $\mathsf{H}_{k^{\prime}}(\mathbf{r})$, which is given by
$\mathsf{V}_{\mathsf{zf},k}(\mathbf{r})=\mathsf{H}_{k}(\mathbf{r})-\frac{\int_{\mathcal{A}}{\mathsf{H}_{k}}(\mathbf{r})\mathsf{H}_{k^{\prime}}^{*}(\mathbf{r})\mathrm{d}\mathbf{r}}{\int_{\mathcal{A}}{\left|\mathsf{H}_{k^{\prime}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}\mathsf{H}_{k^{\prime}}(\mathbf{r}).$ (36)
Substituting (36) into $\mathsf{R}_{\mathsf{ul},k}^{\mathsf{zf}}=\log_{2}(1+\gamma_{\mathsf{ul},k}^{\mathsf{zf}})$ with
and performing some simple manipulations, we can obtain the sum-rate achieved by the ZF detector as follows:
$\mathsf{R}_{\mathsf{ul}}^{\mathsf{zf}}=\sum\nolimits_{k=1}^{2}{\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{ul},k}\mathsf{g}_{k}\big{(}1-\left|\rho\right|^{2}\big{)}\Big{)}}.$ (37)
From the results, it can be observed that $\mathsf{R}_{\mathsf{ul}}^{\mathsf{zf}}\leq\mathsf{C}_{\mathsf{ul}}$, and the gap between them is further illustrated in the numerical results presented in Section VI.
IV Downlink Capacity
In this section, we characterize the channel capacity for CAPA-based downlink communications.
IV-A Single-User Case
For $K=1$, (7b) simplifies to
where ${\mathbbmss{E}}\{{\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}^{*}({\mathbf{s}}^{\prime})\}={\sigma}_{\mathsf{u}}^{2}\delta(\mathbf{s}-{\mathbf{s}}^{\prime})$. Applying a detector ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$ to ${\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s})$ yields
Recalling that the variations of the signal along ${{\mathcal{A}}_{\mathsf{u}}}$ are negligible, we have $\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\approx{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}_{\mathsf{u}})A_{\mathsf{u}}$ and $\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}}\approx\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}}$, which simplifies (40) as follows:
$\gamma_{\mathsf{dl}}\approx\frac{A_{\mathsf{u}}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}}.$ (41)
We note that $\gamma_{\mathsf{dl}}$ does not rely on ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$. Moreover, it is observed from (41) that the SNR is maximized when the current distribution satisfies ${\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\propto{\mathsf{H}}^{*}({\mathbf{r}})$, i.e.,
Remark 3.
Inserting (42) into (41) gives
IV-B Two-User Case
In this part, we focus on the two-user scenario.
@@ IV-B1 Optimal Multiuser Precoding
where ${\mathsf{v}}_{\varepsilon_{k}}=\int_{\mathcal{A}_{\varepsilon_{k}}}{\lvert\mathsf{V}_{{\mathsf{dl}},\varepsilon_{k}}(\mathbf{s})\rvert^{2}}\mathrm{d}\mathbf{s}$ and $\overline{\mathsf{v}}_{\varepsilon_{k}}=\lvert\int_{\mathcal{A}_{\varepsilon_{k}}}{\mathsf{V}}_{{\mathsf{dl}},\varepsilon_{k}}^{*}(\mathbf{s})\mathrm{d}\mathbf{s}\rvert^{2}$. Since the variations of the signal along each user’s aperture is negligible, we have
@@
where $\mathsf{P}$ reflects the power budget. Following the same derivation approach, the sum-rate achieved by the DPC order $1\rightarrow 2$ can also be derived. For downlink Gaussian channels, the maximal DPC sum-rate is independent of the DPC encoding order, and thus $\mathsf{C}_{\mathsf{dl}}^{2\rightarrow 1}$ is essentially the sum-rate capacity [52].
@@
$\begin{split}{\mathsf{C}}_{\mathsf{du}}=\max_{{\mathsf{P}}_{1},{\mathsf{P}}_{2}\geq 0,{\mathsf{P}}_{1}+{\mathsf{P}}_{2}\leq{\mathsf{P}}}\log_{2}(1&+\overline{\gamma}_{\mathsf{du},1}\mathsf{g}_{1}+\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{2}\\ &+\overline{\gamma}_{\mathsf{du},1}\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{1}\mathsf{g}_{2}\overline{\rho}),\end{split}$ (51)
@@
[…]
@@
TABLE II: DPC for downlink CAPA communications.
@@ IV-B2 Capacity Characterization
As proved in Appendix -D, given the power allocation policy $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$, the achievable rates of user $1$ and user $2$ under the DPC order $2\rightarrow 1$ are given by
@@
$\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1}$ $=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}\right),$ (56)
respectively, where $\overline{\gamma}_{\mathsf{dl},k}=\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}{\mathsf{P}}_{k}^{\star}$ for $k\in\{1,2\}$. By comparing the above two expressions with the results in Corollary 1, the following observations are made.
@@
The rates achieved by the DPC order $1\rightarrow 2$ can be derived with similar steps, which are omitted for brevity. In the sequel, we derive the downlink-to-uplink transformation.
@@
=$=\begin{cases}\log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},1}(\mathsf{P})\mathsf{g}_{1}\right)&\xi\geq{\mathsf{P}}\\ \log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},2}(\mathsf{P})\mathsf{g}_{2}\right)&\xi\leq-{\mathsf{P}}\\ \log_{2}\left(1+\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{1}\varepsilon_{2}\overline{\rho}\right)&\mathsf{else}\\ \end{cases},$ (59b)
where $\widetilde{\gamma}_{\mathsf{dl},k}(x)\triangleq\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}x$ for $k\in\{1,2\}$, $\varepsilon_{1}=\widetilde{\gamma}_{\mathsf{dl},1}(\frac{{\mathsf{P}}-\xi}{2})\mathsf{g}_{1}$, and $\varepsilon_{2}=\widetilde{\gamma}_{\mathsf{dl},2}(\frac{{\mathsf{P}}+\xi}{2})\mathsf{g}_{2}$.
@@
According to the uplink-downlink duality, given a power allocation scheme $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$ with ${\mathsf{P}}_{1}+{\mathsf{P}}_{2}={\mathsf{P}}$, we can obtain the capacity region of a dual uplink channel. The downlink capacity region is the convex hull of the union of all these dual uplink capacity regions, which is illustrated in Section VI.
IV-B3 Sum-Rate Achieved by Zero-Forcing Precoder
Although DPC is a potent capacity-achieving scheme, its practical implementation proves challenging. Consequently, non-DPC linear ZF precoding schemes hold practical significance due to the lower complexity. In the following analysis, we aim to evaluate the sum-rate performance achieved by the ZF precoder.
Let $\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})$ denote the ZF source current for user $k$ with $k\in\{1,2\}$. In order to nullify the IUI, i.e., $\int_{{\mathcal{A}}}\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})\mathsf{H}_{k^{\prime}}(\mathbf{r}){\rm{d}}{\mathbf{r}}=0$ for $k^{\prime}\neq k$, and maximize the channel gain of user $k$, $\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})$ is designed as the projection of $\mathsf{H}^{*}_{k}(\mathbf{r})$ onto the space orthogonal to $\mathsf{H}^{*}_{k^{\prime}}(\mathbf{r})$, which is given by
$\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})={\mathsf{H}_{k}^{*}(\mathbf{r})}-\frac{\int_{\mathcal{A}}{\mathsf{H}_{k}^{*}}(\mathbf{r})\mathsf{H}_{k^{\prime}}(\mathbf{r})\mathrm{d}\mathbf{r}}{\int_{\mathcal{A}}{\left|\mathsf{H}_{k^{\prime}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}\mathsf{H}_{k^{\prime}}^{*}(\mathbf{r}).$ (60)
By inserting (60) into the following per-user rate expression
$\mathsf{R}_{\mathsf{dl},k}^{\mathsf{zf}}=\log_{2}\left(1+\frac{|\int_{\mathcal{A}}{}\mathsf{H}_{k}(\mathbf{r})\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})\mathrm{d}\mathbf{r}|^{2}}{\frac{\sigma_{k}^{2}}{A_{\mathsf{u},k}}+|\int_{\mathcal{A}}{}\mathsf{H}_{k}(\mathbf{r})\mathsf{J}_{\mathsf{zf},k^{\prime}}(\mathbf{r})\mathrm{d}\mathbf{r}|^{2}}\right),$ (61)
the sum-rate achieved by the ZF precoding is given as follows:
$\mathsf{R}_{\mathsf{dl}}^{\mathsf{zf}}=\sum\nolimits_{k=1}^{2}{\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{dl},k}^{\mathsf{zf}}\mathsf{g}_{k}\big{(}1-\left|\rho\right|^{2}\big{)}\Big{)}},$ (62)
where $\overline{\gamma}_{\mathsf{dl},k}^{\mathsf{zf}}=\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}{\mathsf{P}}_{k}^{\mathsf{zf}}$ with ${\mathsf{P}}_{k}^{\mathsf{zf}}=\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{zf}},k}(\mathbf{r})\rvert^{2}}\mathrm{d}\mathbf{r}$ being the optimally allocated power that can be obtained by the water-filling algorithm [2, 3].
V Case Studies
As discussed in Remark 2 and Remark 8, the sum-rate capacity expressions derived for both uplink and downlink CAPA communications are applicable to apertures of arbitrary location, shape, and size, among other factors. In this section, we specialize ${\mathcal{A}}$ to several specific structures to unveil more system insights under the two-user scenario.